Find the present value of a series of investment of  $1  at the beginning of each year for 15 years with effective interest rate of 13.5% p.a.


1
Expert's answer
2021-05-11T14:26:26-0400

Solution:

Given, "p=\$1,r=13.5\%\ p.a.=0.135, n=15\ years"

So, present value,"PV=p[\dfrac{1-(1+r)^{-n}}{r}]"

"=1[\dfrac{1-(1+0.135)^{-15}}{0.135}]=\$6.298926\approx\$6.3"